After the vacuum can is heated, I aligned the north cavity again today to about 60% mode matching. however, I'm seeing a weird oscillation in the mixer channel of FSS board which is very big (5 Vpp) and is happening alongside a large power fluctuation in transmitted light. See attached plot from oscilloscope of reflected DC level and the micer channel from FSS. The repition rate of the oscillations make this a 7.35 kHz noise. I have never seen this before. Further, these oscillations are completely unaffected from chaning COM gain or the Fast gain in the FSS path. Comments required.
Code and Data